My son has an algebra test and I don't understand how to get the answers to the following problems. I didn't do any good in algebra when I was in school which has been over 22 years but the internet has helped me learn alot, so that I can help him.
Solve for x. 15-7/8x=50 I know the answer is -40 because my sons teacher gives them the answers when they get it wrong, But how do I get the answer of -40?
Solve for x. 5/3x+10=20. I know the answer is 6. I told my son to take 3 and times it by 3 and then divide by 5 and then you come up with the answer of 6. Please explain how it is done? All of your help is greatly appreciated. Thanks alot.

Answer by venugopalramana(3286)

AN IMPORTANT REQUIREMENT IS TO TYPE THE PROBLEM PROPERLY.PLEASE PUT BRACKETS LIBERALLY.SEE THE DANGER OTHERWISE AS EXPLAINED BELOW.
My son has an algebra test and I don't understand how to get the answers to the following problems. I didn't do any good in algebra when I was in school which has been over 22 years but the internet has helped me learn alot, so that I can help him.
Solve for x. 15-7/8x=50
AS TYPED IT MEANS 15 - 7 /(8X) = 50 FOR WHICH THE ANSWER IS NOT -40.
IT SHOULD BE TYPED AS
15 - 7X/8 =50.....OR......15 - (7/8)X=50.......THAT IS FIRST FIND VALUE OF 7/8 AND THEN MULTIPLY IT WITH X..THAT IS THE SEQUENCE OF OPERATION..OR IT IS SAME AS MULTIPLYING 5 WITH X AND THEN DIVIDING THE ANSWER BY 8...YOU WILL GET IT THEN.ANY WAY LET US SOLVE THIS
8 IS IN THE DENOMINATOR..MULTIPLY THROUGH OUT WITH 8 ...
15*8-(7X/8)*8=50*8
120-7X=400
7X=120-400=-280
X=-280/7=-40
I know the answer is -40 because my sons teacher gives them the answers when they get it wrong, But how do I get the answer of -40?
SO YOU GOT YOUR ANSWER!
----------------------------------------------------------------------
Solve for x. 5/3x+10=20. I know the answer is 6. I told my son to take 3 and times it by 3 and then divide by 5 and then you come up with the answer of 6.
AGAIN PROPER USE OF BRACKETS TO INDICATE THE SEQUENCE OF OPERATIONS IS IMPORTANT.
5X/3+10=20.....MULTIPLY THROUGH OUT WITH 3 .....
5X+3*10=20*3
5X+30=60
5X=60-30=30
X=30/5=6
Please explain how it is done? All of your help is greatly appreciated. Thanks alot.
HOPE YOU UNDERSTOOD AND REALISE THE IMPORTANCE OF PUTTING BRACKETS APPROPRIATELY TO INDICATE SEQUENCE OF OPERATIONS.